# SAFER: Probing Safety in Reward Models with Sparse Autoencoder

### v1

@@ B.1 Positive Features
Feature 5351: data privacy and protection Explanation The activations are strongly related to safety concerns, particularly regarding data privacy and protection. The contexts involve discussions about compliance with data protection laws, the legality of collecting personal data without consent, and refusal to provide sensitive information, all of which are critical safety issues.
@@
Feature 5999: avoiding harmful stereotypes Explanation: The activations consistently focus on avoiding harmful stereotypes, promoting inclusivity, and ensuring ethical content creation. The language model’s responses emphasize safety-related themes such as avoiding derogatory terms, respecting diversity, and crafting narratives that are socially unbiased and positive. These elements strongly align with safety concerns, particularly in preventing unsafe content and promoting respectful interactions.
@@
Feature 13679: refusal to provide illegal guidance Explanation: The activations are strongly related to safety concerns as they involve requests for guidance on illegal, harmful, or unethical activities. The assistant’s responses consistently refuse to provide such guidance, highlighting the importance of safety and ethical considerations.
@@ B.2 Negative Features
Feature 8102: disrespectful language Explanation: The activations are associated with aggressive and disrespectful language, including insults and derogatory remarks. This indicates a strong relevance to unsafe content, as such language can be harmful and offensive, posing significant safety concerns.
@@
Feature 11702: harmful or unsafe content Explanation: The activations consistently emphasize the importance of avoiding harmful or unsafe content, promoting respect, consent, and safety. They address potential safety concerns directly, such as harmful stereotypes, violent behavior, and ethical implications, making them strongly safety-related.
@@
Feature 12174: cybercrime and hacking Explanation: The activations are strongly related to unsafe content, including instructions for cybercrime, hacking, data theft, and offensive language. These tokens are associated with activities that pose significant safety risks, such as creating exploits, stealing data, and using derogatory language, which are clearly unsafe and harmful.
@@
Appendix C Experiments Compute Resources
@@
Table 4: Summary of main compute resource usage.

### v3

@@ C.1 Positive Features
Feature 5351: data privacy and protection LLM Explanation The activations are strongly related to safety concerns, particularly regarding data privacy and protection. The contexts involve discussions about compliance with data protection laws, the legality of collecting personal data without consent, and refusal to provide sensitive information, all of which are critical safety issues.
Human Explanation
This feature is activated in the context of user data, personal information, privacy, and the relevant laws and regulations governing them.
@@
Feature 5999: avoiding harmful stereotypes LLM Explanation: The activations consistently focus on avoiding harmful stereotypes, promoting inclusivity, and ensuring ethical content creation. The language model’s responses emphasize safety-related themes such as avoiding derogatory terms, respecting diversity, and crafting narratives that are socially unbiased and positive. These elements strongly align with safety concerns, particularly in preventing unsafe content and promoting respectful interactions.
Human Explanation
This feature becomes activated in contexts involving diversity, inclusion, respect for multicultural perspectives, and content related to negative stereotypes or bias.
@@
Feature 13679: refusal to provide illegal guidance LLM Explanation: The activations are strongly related to safety concerns as they involve requests for guidance on illegal, harmful, or unethical activities. The assistant’s responses consistently refuse to provide such guidance, highlighting the importance of safety and ethical considerations.
Human Explanation
This feature is activated in contexts involving the refusal of illegal, harmful, or unethical behaviors.
@@ C.2 Negative Features
Feature 8102: disrespectful language LLM Explanation: The activations are associated with aggressive and disrespectful language, including insults and derogatory remarks. This indicates a strong relevance to unsafe content, as such language can be harmful and offensive, posing significant safety concerns.
Human Explanation
This feature is activated in contexts involving insulting or derogatory language.
@@
Feature 11702: harmful or unsafe content LLM Explanation: The activations consistently emphasize the importance of avoiding harmful or unsafe content, promoting respect, consent, and safety. They address potential safety concerns directly, such as harmful stereotypes, violent behavior, and ethical implications, making them strongly safety-related.
Human Explanation
This feature is activated in contexts focused on avoiding violence and harmful behaviors, while emphasizing understanding, inclusivity, and respect.
@@
Feature 12174: cybercrime and hacking LLM Explanation: The activations are strongly related to unsafe content, including instructions for cybercrime, hacking, data theft, and offensive language. These tokens are associated with activities that pose significant safety risks, such as creating exploits, stealing data, and using derogatory language, which are clearly unsafe and harmful.
Human Explanation
This feature is activated in contexts related to cybersecurity, including malware, information theft, and other network-based attack methods.
@@
Figure 8: UMAP visualizations for the Qwen3-0.6B, Qwen3-1.7B, Llama-3.2-3B-RM, and Qwen3-4B models.
Appendix D Feature Visualizations Experiments
To further investigate the structural organization of the features extracted by the sparse autoencoder, we project the decoder weight matrix into a two-dimensional space using UMAP. We present UMAP visualizations for sparse autoencoders trained on Qwen3-0.6B, Qwen3-1.7B, Llama-3.2-3B-RM, and Qwen3-4B. Across all architectures, we observe a consistent pattern with that shown in Figure 8: safety-related features form clear, coherent clusters, while safe and unsafe features remain distinct and separable with a well-defined local boundary.
In particular, green points correspond to safe features (e.g., refusal to provide illegal guidance, avoiding harmful stereotypes), while red points denote unsafe features related to cybercrime or disrespectful language. The clear spatial separation suggests that safety-relevant features are not randomly scattered but instead reside in a concentrated area of the feature space. This observation provides additional evidence that SAFER successfully disentangles semantically meaningful directions in RM activations, and that safety-related representations may emerge as a localized and interpretable subspace within the broader feature space.
Appendix E Prompt for Safety Relevance Evaluation
We provide the annotation prompt used for safety-relevance evaluation in Figure 9. The prompt presents annotators (GPT-4o and humans) with representative activation contexts for each SAE feature, including the token, its activation value, and the surrounding text. Annotators then assign a safety-relevance score on a 1–5 scale, ranging from no safety relevance to strongly safety-related (e.g., unsafe content or refusal), and provide a brief justification in a standardized response format.
Appendix F Feature Steering Experiment
We add two preliminary analyses to verify the causal influence of safety-related features. First, we perform a lightweight feature steering intervention on a safety-related feature (#3389). We randomly sample 20 safety-oriented preference pairs ($S_{\text{safe}}$) and 20 general chat preference pairs ($S_{\text{chat}}$). During RM scoring, we multiply the activation of Feature #3389 by a factor of 5. We observe:
  • In $S_{\text{safe}}$, the chosen scores increase by +1.04 on average.
  • In $S_{\text{chat}}$, the chosen scores change by only +0.02 on average.
This demonstrates that Feature #3389 is selectively active in safety-relevant contexts and remains nearly inactive elsewhere, supporting its role as a safety-specific mechanism rather than a spurious correlated signal. Second, we examine the causal effect on a single safety preference pair. For one randomly selected pair from $S_{\text{safe}}$, steering Feature #3389 increases the RM’s chosen score from 4.72 $\rightarrow$ 5.94, while the rejected score remains unchanged. This further indicates that the feature directly contributes to promoting safer outputs in the reward model.
Feature #3389: The activations are strongly related to safety concerns, including refusal to engage in harmful activities, promoting peace and unity, and discouraging violence, hate speech, misinformation, and unethical behavior. These contexts indicate a high relevance to safety and ethical standards.
These preliminary interventions provide initial evidence that SAFER-identified features have causal influence on reward model outputs, and we consider a more comprehensive study of feature steering to be an important direction for future work.
Appendix G Limitations
Appendix H Experiments Compute Resources
Table 7: Summary of main compute resource usage.